# Couples can be tractable: New algorithms and hardness results for the Hospitals / Residents problem with Couples

### v2

@@ 4.1 Sub-responsive and sub-complete preferences
In this section we show that hrc remains NP-hard even with sub-responsive and sub-complete couple preferences. We provide a reduction from com smti, which is the problem of deciding, given an instance of Stable Marriage with Ties and Incomplete lists, whether a complete stable matching (i.e., a stable matching that matches all agents) exists (this problem is defined formally in [22, Section 1.3.5], for example).
@@
hrc is NP-hard even if each couple has sub-responsive and sub-complete preferences. This holds even with unit hospital capacities, and even if each preference list is of length at most 4. This also holds if either (i) the couples are half-separable, or (ii) every couple is separable or of type-a. In cases (i) and (ii), each preference list is of length at most 6 and 8, respectively.
@@
Let $U$ be the set of men in $I$ and let $W$ be the set of women in $I$. Let $G=(U\cup W,E)$ denote the underlying bipartite graph of $I$, and let $m_{I}$ denote $|E|$. Let $W^{s}$ denote the set of women in $W$ whose preference lists are strictly ordered in $I$, and let $W^{t}=W\setminus W^{s}$. Then every woman in $W^{t}$ has a preference list in $I$ that is a tie of length 2.
@@
In the case that we want the couples to be also half-separable, we allow the doctors $c_{u}^{1},c_{u}^{3},c_{u}^{5}$ and $c_{w}^{1}$ to be matched alone (but not their partners). Also, in this case the couples’ preferences lists in $I^{\prime}$ are extended as shown in Figure 5. It is straightforward to verify that the preference list of each single doctor is of length at most 4, the preference list of each couple is of length at most 6, and the preference list of each hospital is of length at most 3.
Alternatively, if we want the couples to be separable and of type-a, a similar modification to the preference lists of the couples $(c_{w}^{1},c_{w}^{2})$ to that shown in Figure 5 can be used to make them separable. All other couples are of type-a. We then obtain that the preference list of each single doctor is of length at most 4, the preference list of each couple is of length at most 8, and the preference list of each hospital is of length at most 3.
$\begin{array}[]{rll}(c_{u}^{1},c_{u}^{2}):&(f_{u},h_{u}^{2})\succ(h_{u}^{1},h_{u}^{2})\succ(f_{u},\emptyset)\succ(h_{u}^{1},\emptyset)&(u\in U)\\ (c_{u}^{3},c_{u}^{4}):&(h_{u}^{2},h_{u}^{3})\succ(h_{u}^{2},\emptyset)&(u\in U)\\ (c_{u}^{5},c_{u}^{6}):&(h_{u}^{3},h_{u}^{1})\succ(h_{u}^{3},\emptyset)&(u\in U)\\ (c_{w}^{1},c_{w}^{2}):&(h_{w}^{2},h_{w}^{2})\succ(h_{w}^{1},h_{w}^{2})\succ(h_{w}^{2},h_{w}^{3})\succ(h_{w}^{1},h_{w}^{3})\succ(h_{w}^{2},\emptyset)\succ(h_{w}^{1},\emptyset)&(w\in W)\\ \end{array}$
Figure 5: Extended preferences for the half-separable case
@@
Let $M$ be a complete stable matching in $I$. We create a matching $M^{\prime}$ in $I^{\prime}$ as follows. For each $(u,w)\in M$ we add the pair $(d_{u},h_{w})$ to $M^{\prime}$ if $w\in W^{s}$, and the pair $(d_{u},h_{w}^{1})$ or $(d_{u},h_{w}^{3})$ to $M^{\prime}$ if $w\in W^{t}$ (depending on which of these two hospitals is acceptable to $d_{u}$). Then, for each $w\in W^{t}$, if $h_{w}^{1}$ has a single doctor $d_{u}$, we match the couple $(c_{w}^{1},c_{w}^{2})$ to $(h_{w}^{2},h_{w}^{3})$, otherwise we match the couple to $(h_{w}^{1},h_{w}^{2})$. Finally, for each $u\in U$, we match $(c_{u}^{1},c_{u}^{2})$ to $(f_{u},h_{u}^{2})$, and $(c_{u}^{5},c_{u}^{6})$ to $(h_{u}^{3},h_{u}^{1})$.
@@
Hence, no doctor or couple can block $M^{\prime}$ in $I^{\prime}$, so $M^{\prime}$ is stable in $I^{\prime}$. By observing that $c_{3}^{u}$ $(u\in U)$ cannot block $M^{\prime}$ alone either, and every other couple is matched in $M^{\prime}$, it follows that $M^{\prime}$ is stable too in $I^{\prime}$ in the half-separable case. ∎
@@
Proof. Let $M^{\prime}$ be a stable matching in $I^{\prime}$. First we show that each doctor $d_{u}$ is matched to some $h_{w}$ or $h_{w}^{j}$ hospital. Suppose this is not the case for some doctor $d_{u}$. Then, as $d_{u}$ is the first choice of $f_{u}$, it follows that $(d_{u},f_{u})\in M^{\prime}$. But then there is no stable allocation of $(c_{u}^{1},c_{u}^{2}),(c_{u}^{3},c_{u}^{4}),(c_{u}^{5},c_{u}^{6})$ to $h_{u}^{1},h_{u}^{2},h_{u}^{3}$. Indeed, if the couples are not half-separable, then it is easy to verify that each of the three possible matchings is blocked by a couple. If they are half-separable, then without loss of generality suppose that $c_{u}^{1}$ is matched in $M^{\prime}$ but $c_{u}^{2}$ is not. Then $(c_{u}^{1},c_{u}^{2})$ blocks $M^{\prime}$ in $I^{\prime}$, a contradiction.
Next, we show that there is no woman $w\in W$ such that the hospital(s) corresponding to $w$ receive more then one doctor $d_{u}$ in $M^{\prime}$. For a woman in $W^{s}$, this is trivial. Let $w\in W^{t}$ and suppose to the contrary that both $h_{w}^{1}$ and $h_{w}^{3}$ are assigned a single doctor in $M^{\prime}$. If the couples are not half-separable, then $(c_{w}^{1},c_{w}^{2})$ cannot be matched, so the couple blocks $M^{\prime}$ in $I^{\prime}$ with $(h_{w}^{2},h_{w}^{3})$, contradiction. If the couples are half-separable, then only $c_{w}^{1}$ can be matched, but $(c_{w}^{1},c_{w}^{2})$ still blocks $M^{\prime}$ in $I^{\prime}$ with $(h_{w}^{2},h_{w}^{3})$.
@@
Remark 21.
We remark that the last entries $(h_{w}^{1},h_{w}^{3})$ in the preference lists of the couples $(c_{w}^{1},c_{w}^{2})$ are redundant, and they are only used to satisfy sub-completeness. So, by dropping them we obtain that hrc is NP-hard, even if the couples are sub-responsive, half-separable and for each couple $C$, there is a hospital $h$ that appears in each entry in $\succ_{C}$.
@@ 4.2 Dual markets
In this section we present our hardness results for hrc-dual market, even in the case when both sides have master lists. We begin by showing that, given an instance of hrc-dual market, it is NP-complete to decide whether a complete stable matching (i.e., a stable matching in which all single doctors and couples are matched) exists,
@@
In the joint preference list of a couple $(x_{4i+r},k_{4i+r})(x_{4i+r}\in X,k_{4i+r}\in K)$, if $r\in\{0,1\}$, the symbol $c(x_{4i+r})$ denotes the hospital $c_{j}^{s}\in C_{B}^{\prime}$ such that the $(r+1)$th occurrence of literal $v_{i}$ appears at position $s$ of $c_{j}$. Similarly if $r\in\{2,3\}$ then the symbol $c(x_{4i+r})$ denotes the hospital $c_{j}^{s}\in C_{B}^{\prime}$ such that the $(r-1)$th occurrence of literal $\bar{v_{i}}$ appears at position $s$ of $c_{j}$. Also in the preference list of a hospital $c_{j}^{s}\in C_{B}^{\prime}$, if literal $v_{i}$ appears at position $s$ of clause $c_{j}\in C_{B}$, the symbol $x(c_{j}^{s})$ denotes the doctor $x_{4i+r-1}$ where $r=1,2$ according as this is the first or second occurrence of literal $v_{i}$ in $B$. Otherwise if literal $\bar{v_{i}}$ appears at position $s$ of clause $c_{j}\in C_{B}$, the symbol $x(c_{j}^{s})$ denotes the doctor $x_{4i+r+1}$ where $r=1,2$ according as this is the first or second occurrence of literal $\bar{v_{i}}$ in $B$.
@@
The preference lists of the doctors and hospitals in $I$ are shown in Figure 6. Clearly each preference list is of length at most 3. The doctors in $P_{j}\cup\{c_{j}^{s}:1\leq s\leq 3\}\cup\{z_{j}^{s}:1\leq s\leq 5\}\cup\{s_{j},t_{j}\}$ constitute a clause gadget for each $j$ ($1\leq j\leq m$) and such a gadget is illustrate in Figure 7. Similarly the doctors in $X_{i}\cup K_{i}$ and the hospitals in $Y_{i}\cup L_{i}$ constitute a variable gadget for each $i$ ($0\leq i\leq n-1)$ and such a gadget is illustrated in Figure 8. In each of Figures 7 and 8, two doctors involved in a couple are shown using dashed lines. Numbers beside full lines indicate preference list rankings.
$\begin{array}[]{rll}(x_{4i},k_{4i}):&(y_{4i},l_{4i})~{}~{}(c(x_{4i}),l_{4i+1})~{}~{}(y_{4i+1},l_{4i+1})&(0\leq i\leq n-1)\\ (x_{4i+1},k_{4i+1}):&(y_{4i+1},l_{4i+1})~{}~{}(c(x_{4i+1}),l_{4i+2})~{}~{}(y_{4i+2},l_{4i+2})&(0\leq i\leq n-1)\\ (x_{4i+2},k_{4i+2}):&(y_{4i+3},l_{4i+3})~{}~{}(c(x_{4i+2}),l_{4i+2})~{}~{}(y_{4i+2},l_{4i+2})&(0\leq i\leq n-1)\\ (x_{4i+3},k_{4i+3}):&(y_{4i},l_{4i})~{}~{}(c(x_{4i+3}),l_{4i+3})~{}~{}(y_{4i+3},l_{4i+3})&(0\leq i\leq n-1)\\ (p_{j}^{1},p_{j}^{4}):&(z^{1}_{j},z^{2}_{j})~{}~{}(c_{j}^{1},z^{3}_{j})&(1\leq j\leq m)\vspace{1mm}\\ (p_{j}^{2},p_{j}^{5}):&(z^{1}_{j},z^{2}_{j})~{}~{}(c_{j}^{2},z^{4}_{j})&(1\leq j\leq m)\vspace{1mm}\\ (p_{j}^{3},p_{j}^{6}):&(z^{1}_{j},z^{2}_{j})~{}~{}(c_{j}^{3},z^{5}_{j})&(1\leq j\leq m)\vspace{1mm}\\ s_{j}:&c_{j}^{1}~{}~{}c_{j}^{2}~{}~{}c_{j}^{3}&(1\leq j\leq m)\\ t_{j}:&z^{3}_{j}~{}~{}z^{4}_{j}~{}~{}z^{5}_{j}&(1\leq j\leq m)\vspace{1mm}\\ \\ y_{4i}:&x_{4i}~{}~{}x_{4i+3}&(0\leq i\leq n-1)\\ y_{4i+1}:&x_{4i+1}~{}~{}x_{4i}&(0\leq i\leq n-1)\\ y_{4i+2}:&x_{4i+1}~{}~{}x_{4i+2}&(0\leq i\leq n-1)\\ y_{4i+3}:&x_{4i+2}~{}~{}x_{4i+3}&(0\leq i\leq n-1)\\ l_{4i}:&k_{4i+3}~{}~{}k_{4i}&(0\leq i\leq n-1)\\ l_{4i+1}:&k_{4i}~{}~{}k_{4i+1}&(0\leq i\leq n-1)\\ l_{4i+2}:&k_{4i+2}~{}~{}k_{4i+1}&(0\leq i\leq n-1)\\ l_{4i+3}:&k_{4i+3}~{}~{}k_{4i+2}&(0\leq i\leq n-1)\\ z^{1}_{j}:&p_{j}^{1}~{}~{}p_{j}^{2}~{}~{}p_{j}^{3}&(1\leq j\leq m)\vspace{1mm}\\ z^{2}_{j}:&p_{j}^{6}~{}~{}p_{j}^{5}~{}~{}p_{j}^{4}&(1\leq j\leq m)\vspace{1mm}\\ z^{2+s}_{j}:&p_{j}^{3+s}~{}~{}t_{j}&(1\leq j\leq m\wedge 1\leq s\leq 3)\vspace{1mm}\\ c_{j}^{s}:&p_{j}^{s}~{}~{}x(c_{j}^{s})~{}~{}s_{j}&(1\leq j\leq m\wedge 1\leq s\leq 3)\vspace{1mm}\\ \end{array}$
@@
Let $j$ $(1\leq j\leq m)$ be given. Then $c_{j}$ contains at least one literal that is true under $f$. Suppose this literal occurs at position $s$ of $c_{j}$ $(1\leq s\leq 3)$. Then add the pairs ($s_{j},c^{s}_{j}$), ($p^{s}_{j},z^{1}_{j}$) and ($p_{j}^{s+3},z^{2}_{j}$) to $M$. For each $b\in\{1,2,3\}\setminus\{s\}$, add the pairs ($p^{b}_{j},c^{b}_{j}$) and ($p_{j}^{b+3},z_{j}^{b+2}$) to $M$. Finally add the pair ($t_{j},z_{j}^{s+2}$) to $M$.
@@
Suppose that $(x_{4i+r},k_{4i+r})$ blocks $M$ with hospitals in $Y\cup L$ for some $i$ ($10\leq i\leq n-$ 1) and $r$ ($0\leq r\leq 3$). If $T_{i}\subseteq M$ then $r\in\{2,3\}$ so that $(x_{4i+r},k_{4i+r})$ have their third choice. But then each of $y_{4i}$ and $l_{4i+3}$ have their first choice, so $(x_{4i+r},k_{4i+r})$ cannot block $M$ after all, a contradiction. If $F_{i}\subseteq M$ then $r\in\{0,1\}$ so that $(x_{4i+r},k_{4i+r})$ have their third choice. But then each of $l_{4i}$ and $l_{4i+1}$ have their first choice, so $(x_{4i+r},k_{4i+r})$ cannot block $M$ after all, a contradiction.
@@
Proof. Conversely suppose that $M$ is a complete stable matching in $I$. We form a truth assignment $f$ in $B$ as follows. For each $i$ ($10\leq i\leq n-$ 1), $M\cap((X_{i}\times Y_{i})\cup(K_{i}\times L_{i}))$ is a perfect matching of ($X_{i}\cup Y_{i})\cup(K_{i}\cup L_{i}$). If $M\cap((X_{i}\times Y_{i})\cup(K_{i}\times L_{i}))=T_{i}$, set $v_{i}$ to be true under $f$. Otherwise $M\cap((X_{i}\times Y_{i})\cup(K_{i}\times L_{i}))=F_{i}$, in which case we set $v_{i}$ to be false under $f$.
@@
Now let $c_{j}$ be a clause in $C_{B}$ ($1\leq j\leq m$). There exists some $s$ ($1\leq s\leq 3$) such that $(s_{j},c_{j}^{s})\in M$. Let $x_{4i+r}=x(c_{j}^{s})$, for some $i$ ($10\leq i\leq n-$ 1) and $r$ ($0\leq r\leq 3$). If $r\in\{0,1\}$ then $(x_{4i+r},y_{4i+r})\in M$ (or equivalently $(k_{4i+r},l_{4i+r})\in M$) by the stability of $M$. Thus variable $v_{i}$ is true under $f$, and hence clause $c_{j}$ is true under $f$, since literal $v_{i}$ occurs in $c_{j}$. If $r\in\{2,3\}$ then $(x_{4i+r},y_{4i+r+1})\in M$ (or equivalently $(k_{4i+r},l_{4i+r+1})\in M$) (where addition is taken modulo $4$) by the stability of $M$. Thus variable $v_{i}$ is false under $f$, and hence clause $c_{j}$ is true under $f$, since literal $\bar{v_{i}}$ occurs in $c_{j}$. Hence $f$ is a satisfying truth assignment of $B$. ∎  $\square$
@@
Let $I$ be the instance of hrc-dual market as constructed in the proof of Theorem 22 from an instance of (2,2)-e3-sat. We now add additional doctors and hospitals to $I$, in the form of a series of enforcer gadgets, to obtain a new instance $I^{\prime}$ of hrc-dual market as follows, with the aim that every stable matching in $I^{\prime}$ must be complete. For each $i$ ($10\leq i\leq n-$ 1) and $r$ ($0\leq r\leq 3$), we add an enforcer gadget to $y_{4i+r}$ comprising two new couples, $(u_{4i+r}^{1},u_{4i+r}^{2})$ and $(u_{4i+r}^{3},u_{4i+r}^{4})$, a single doctor $u_{4i+r}^{5}$, and four new hospitals $h_{4i+r}^{s}$ ($1\leq s\leq 4$), all of capacity 1. The preference lists of the newly added agents corresponding to $y_{4i+r}\in Y$ are shown in Figure 9 and the enforcer gadget is illustrated in Figure 10. In the preference list of $y_{4i+r}$ in $I^{\prime}$, $P(y_{4i+r})$ represents the preference list of $y_{4i+r}$ in $I$.
$\begin{array}[]{rll}(u^{1}_{4i+r},u^{2}_{4i+r}):&(h^{1}_{4i+r},h^{2}_{4i+r}){}{}{}{}&(0\leq i\leq n-1,0\leq r\leq 3)\\ (u^{3}_{4i+r},u^{4}_{4i+r}):&(h^{1}_{4i+r},h^{4}_{4i+r})~{}~{}(h^{3}_{4i+r},h^{2}_{4i+r}){}{}&(0\leq i\leq n-1,0\leq r\leq 3)\\ u^{5}_{4i+r}:&y_{4i+r}~{}~{}h^{1}_{4i+r}{}{}&(0\leq i\leq n-1,0\leq r\leq 3)\\ \\ y_{4i+r}:&P(y_{4i+r})~{}~{}u^{5}_{4i+r}&(0\leq i\leq n-1,0\leq r\leq 3)\vspace{1mm}\\ h^{1}_{4i+r}:&u^{5}_{4i+r}~{}~{}u^{1}_{4i+r}~{}~{}u^{3}_{4i+r}&(0\leq i\leq n-1,0\leq r\leq 3)\vspace{1mm}\\ h^{2}_{4i+r}:&u^{4}_{4i+r}~{}~{}u^{2}_{4i+r}&(0\leq i\leq n-1,0\leq r\leq 3)\vspace{1mm}\\ h^{3}_{4i+r}:&u^{3}_{4i+r}&(0\leq i\leq n-1,0\leq r\leq 3)\vspace{1mm}\\ h^{4}_{4i+r}:&u^{4}_{4i+r}&(0\leq i\leq n-1,0\leq r\leq 3)\vspace{1mm}\\ \end{array}$

### v3

@@ 4.1 Sub-responsive and sub-complete preferences
In this section we show that hrc remains NP-hard even with sub-responsive and sub-complete couple preferences. We provide a reduction from com smti, which is the problem of deciding, given an instance of Stable Marriage with Ties and Incomplete lists, whether a complete stable matching (i.e., a stable matching that matches all agents) exists (this problem is defined formally in [23, Section 1.3.5], for example).
@@
hrc is NP-hard even if each couple has sub-responsive and sub-complete preferences. This holds even with unit hospital capacities, and even if each preference list is of length at most 4.
@@
Let $U$ be the set of men in $I$ and let $W$ be the set of women in $I$. Let $G=(U\cup W,E)$ denote the underlying bipartite graph of $I$, and let $m_{I}$ denote $|E|$. Let $W^{s}$ denote the set of women in $W$, whose preference lists are strictly ordered in $I$, and let $W^{t}=W\setminus W^{s}$. Then, every woman in $W^{t}$ has a preference list in $I$ that is a tie of length 2.
@@
Let $M$ be a complete stable matching in $I$. We create a matching $M^{\prime}$ in $I^{\prime}$ as follows. For each $(u,w)\in M$ we add the pair $(d_{u},h_{w})$ to $M^{\prime}$, if $w\in W^{s}$, and the pair $(d_{u},h_{w}^{1})$ or $(d_{u},h_{w}^{3})$ to $M^{\prime}$, if $w\in W^{t}$ (depending on which one of these two hospitals is acceptable to $d_{u}$). Then, for each $w\in W^{t}$, if $h_{w}^{1}$ has a single doctor $d_{u}$, we match the couple $(c_{w}^{1},c_{w}^{2})$ to $(h_{w}^{2},h_{w}^{3})$, otherwise we match the couple to $(h_{w}^{1},h_{w}^{2})$. Finally, for each $u\in U$, we match $(c_{u}^{1},c_{u}^{2})$ to $(f_{u},h_{u}^{2})$, and $(c_{u}^{5},c_{u}^{6})$ to $(h_{u}^{3},h_{u}^{1})$.
@@
Hence, no doctor or couple can block $M^{\prime}$ in $I^{\prime}$, so $M^{\prime}$ is stable in $I^{\prime}$. ∎
@@
Proof. Let $M^{\prime}$ be a stable matching in $I^{\prime}$. First we show that each doctor $d_{u}$ is matched to some $h_{w}$ or $h_{w}^{j}$ hospital. Suppose this is not the case for some doctor $d_{u}$. Then, as $d_{u}$ is the first choice of $f_{u}$, it follows that $(d_{u},f_{u})\in M^{\prime}$. But then, there is no stable allocation of $(c_{u}^{1},c_{u}^{2}),(c_{u}^{3},c_{u}^{4}),(c_{u}^{5},c_{u}^{6})$ to $h_{u}^{1},h_{u}^{2},h_{u}^{3}$. Indeed, it is easy to verify that each of the three possible matchings is blocked by a couple.
Next, we show that there is no woman $w\in W$ such that the hospital(s) corresponding to $w$ receive more then one doctor $d_{u}$ in $M^{\prime}$. For a woman in $W^{s}$, this is trivial. Let $w\in W^{t}$ and suppose to the contrary that both $h_{w}^{1}$ and $h_{w}^{3}$ are assigned a single doctor in $M^{\prime}$. Then, $(c_{w}^{1},c_{w}^{2})$ cannot be matched, so the couple blocks $M^{\prime}$ in $I^{\prime}$ with $(h_{w}^{2},h_{w}^{3})$, contradiction.
@@ 4.2 Dual markets
In this section we present our hardness results for hrc-dual market, even in the case when both sides have master lists. We begin by showing that, given an instance of hrc-dual market, it is NP-complete to decide whether a complete stable matching (i.e., a stable matching in which all single doctors and couples are matched) exists.
@@
In the joint preference list of a couple $(x_{4i+r},k_{4i+r})(x_{4i+r}\in X,k_{4i+r}\in K)$, if $r\in\{0,1\}$, the symbol $c(x_{4i+r})$ denotes the hospital $c_{j}^{s}\in C_{B}^{\prime}$ such that the $(r+1)$th occurrence of literal $v_{i}$ appears at position $s$ of $c_{j}$. Similarly, if $r\in\{2,3\}$ then the symbol $c(x_{4i+r})$ denotes the hospital $c_{j}^{s}\in C_{B}^{\prime}$ such that the $(r-1)$th occurrence of literal $\bar{v_{i}}$ appears at position $s$ of $c_{j}$. Also in the preference list of a hospital $c_{j}^{s}\in C_{B}^{\prime}$, if literal $v_{i}$ appears at position $s$ of clause $c_{j}\in C_{B}$, the symbol $x(c_{j}^{s})$ denotes the doctor $x_{4i+r-1}$ where $r=1,2$ according as this is the first or second occurrence of literal $v_{i}$ in $B$. Otherwise if literal $\bar{v_{i}}$ appears at position $s$ of clause $c_{j}\in C_{B}$, the symbol $x(c_{j}^{s})$ denotes the doctor $x_{4i+r+1}$ where $r=1,2$ according as this is the first or second occurrence of literal $\bar{v_{i}}$ in $B$.
@@
The preference lists of the doctors and hospitals in $I$ are shown in Figure 5. Clearly each preference list is of length at most 3. The doctors in $P_{j}\cup\{c_{j}^{s}:1\leq s\leq 3\}\cup\{z_{j}^{s}:1\leq s\leq 5\}\cup\{s_{j},t_{j}\}$ constitute a clause gadget for each $j$ ($1\leq j\leq m$) and such a gadget is illustrate in Figure 6. Similarly the doctors in $X_{i}\cup K_{i}$ and the hospitals in $Y_{i}\cup L_{i}$ constitute a variable gadget for each $i$ ($0\leq i\leq n-1)$ and such a gadget is illustrated in Figure 7. In each of Figures 6 and 7, two doctors involved in a couple are shown using dashed lines. Numbers beside full lines indicate preference list rankings.
$\begin{array}[]{rll}(x_{4i},k_{4i}):&(y_{4i},l_{4i})\succ(c(x_{4i}),l_{4i+1})\succ(y_{4i+1},l_{4i+1})&(0\leq i\leq n-1)\\ (x_{4i+1},k_{4i+1}):&(y_{4i+1},l_{4i+1})\succ(c(x_{4i+1}),l_{4i+2})\succ(y_{4i+2},l_{4i+2})&(0\leq i\leq n-1)\\ (x_{4i+2},k_{4i+2}):&(y_{4i+3},l_{4i+3})\succ(c(x_{4i+2}),l_{4i+2})\succ(y_{4i+2},l_{4i+2})&(0\leq i\leq n-1)\\ (x_{4i+3},k_{4i+3}):&(y_{4i},l_{4i})\succ(c(x_{4i+3}),l_{4i+3})\succ(y_{4i+3},l_{4i+3})&(0\leq i\leq n-1)\\ (p_{j}^{1},p_{j}^{4}):&(z^{1}_{j},z^{2}_{j})\succ(c_{j}^{1},z^{3}_{j})&(1\leq j\leq m)\vspace{1mm}\\ (p_{j}^{2},p_{j}^{5}):&(z^{1}_{j},z^{2}_{j})\succ(c_{j}^{2},z^{4}_{j})&(1\leq j\leq m)\vspace{1mm}\\ (p_{j}^{3},p_{j}^{6}):&(z^{1}_{j},z^{2}_{j})\succ(c_{j}^{3},z^{5}_{j})&(1\leq j\leq m)\vspace{1mm}\\ s_{j}:&c_{j}^{1}\succ c_{j}^{2}\succ c_{j}^{3}&(1\leq j\leq m)\\ t_{j}:&z^{3}_{j}\succ z^{4}_{j}\succ z^{5}_{j}&(1\leq j\leq m)\vspace{1mm}\\ \\ y_{4i}:&x_{4i}\succ x_{4i+3}&(0\leq i\leq n-1)\\ y_{4i+1}:&x_{4i+1}\succ x_{4i}&(0\leq i\leq n-1)\\ y_{4i+2}:&x_{4i+1}\succ x_{4i+2}&(0\leq i\leq n-1)\\ y_{4i+3}:&x_{4i+2}\succ x_{4i+3}&(0\leq i\leq n-1)\\ l_{4i}:&k_{4i+3}\succ k_{4i}&(0\leq i\leq n-1)\\ l_{4i+1}:&k_{4i}\succ k_{4i+1}&(0\leq i\leq n-1)\\ l_{4i+2}:&k_{4i+2}\succ k_{4i+1}&(0\leq i\leq n-1)\\ l_{4i+3}:&k_{4i+3}\succ k_{4i+2}&(0\leq i\leq n-1)\\ z^{1}_{j}:&p_{j}^{1}\succ p_{j}^{2}\succ p_{j}^{3}&(1\leq j\leq m)\vspace{1mm}\\ z^{2}_{j}:&p_{j}^{6}\succ p_{j}^{5}\succ p_{j}^{4}&(1\leq j\leq m)\vspace{1mm}\\ z^{2+s}_{j}:&p_{j}^{3+s}\succ t_{j}&(1\leq j\leq m\wedge 1\leq s\leq 3)\vspace{1mm}\\ c_{j}^{s}:&p_{j}^{s}\succ x(c_{j}^{s})\succ s_{j}&(1\leq j\leq m\wedge 1\leq s\leq 3)\vspace{1mm}\\ \end{array}$
@@
Let $j$ $(1\leq j\leq m)$ be given. Then $c_{j}$ contains at least one literal that is true under $f$. Suppose this literal occurs at position $s$ of $c_{j}$ $(1\leq s\leq 3)$. Then add the pairs ($s_{j},c^{s}_{j}$), ($p^{s}_{j},z^{1}_{j}$) and ($p_{j}^{s+3},z^{2}_{j}$) to $M$. For each $b\in\{1,2,3\}\setminus\{s\}$, add the pairs ($p^{b}_{j},c^{b}_{j}$) and ($p_{j}^{b+3},z_{j}^{b+2}$) to $M$. Finally, add the pair ($t_{j},z_{j}^{s+2}$) to $M$.
@@
Suppose that $(x_{4i+r},k_{4i+r})$ blocks $M$ with hospitals in $Y\cup L$ for some $i$ ($10\leq i\leq n-$ 1) and $r$ ($0\leq r\leq 3$). If $T_{i}\subseteq M$ then $r\in\{2,3\}$ so that $(x_{4i+r},k_{4i+r})$ have their third choice. But then, each of $y_{4i}$ and $l_{4i+3}$ have their first choice, so $(x_{4i+r},k_{4i+r})$ cannot block $M$ after all, a contradiction. If $F_{i}\subseteq M$ then $r\in\{0,1\}$, so $(x_{4i+r},k_{4i+r})$ have their third choice. But then, each of $l_{4i}$ and $l_{4i+1}$ have their first choice, so $(x_{4i+r},k_{4i+r})$ cannot block $M$ after all, a contradiction.
@@
Proof. Conversely, suppose that $M$ is a complete stable matching in $I$. We form a truth assignment $f$ in $B$ as follows. For each $i$ ($10\leq i\leq n-$ 1), $M\cap((X_{i}\times Y_{i})\cup(K_{i}\times L_{i}))$ is a perfect matching of ($X_{i}\cup Y_{i})\cup(K_{i}\cup L_{i}$). If $M\cap((X_{i}\times Y_{i})\cup(K_{i}\times L_{i}))=T_{i}$, set $v_{i}$ to be true under $f$. Otherwise, $M\cap((X_{i}\times Y_{i})\cup(K_{i}\times L_{i}))=F_{i}$, in which case we set $v_{i}$ to be false under $f$.
@@
Now, let $c_{j}$ be a clause in $C_{B}$ ($1\leq j\leq m$). There exists some $s$ ($1\leq s\leq 3$) such that $(s_{j},c_{j}^{s})\in M$. Let $x_{4i+r}=x(c_{j}^{s})$, for some $i$ ($10\leq i\leq n-$ 1) and $r$ ($0\leq r\leq 3$). If $r\in\{0,1\}$, then $(x_{4i+r},y_{4i+r})\in M$ (or equivalently $(k_{4i+r},l_{4i+r})\in M$) by the stability of $M$. Thus, variable $v_{i}$ is true under $f$, and hence clause $c_{j}$ is true under $f$, since literal $v_{i}$ occurs in $c_{j}$. If $r\in\{2,3\}$, then $(x_{4i+r},y_{4i+r+1})\in M$ (or equivalently $(k_{4i+r},l_{4i+r+1})\in M$) (where addition is taken modulo $4$) by the stability of $M$. Thus variable $v_{i}$ is false under $f$, and hence clause $c_{j}$ is true under $f$, since literal $\bar{v_{i}}$ occurs in $c_{j}$. Hence $f$ is a satisfying truth assignment of $B$. ∎  $\square$
@@
Let $I$ be the instance of hrc-dual market as constructed in the proof of Theorem 22 from an instance of (2,2)-e3-sat. We now add additional doctors and hospitals to $I$, in the form of a series of enforcer gadgets, to obtain a new instance $I^{\prime}$ of hrc-dual market as follows, with the aim that every stable matching in $I^{\prime}$ must be complete. For each $i$ ($10\leq i\leq n-$ 1) and $r$ ($0\leq r\leq 3$), we add an enforcer gadget to $y_{4i+r}$ comprising two new couples, $(u_{4i+r}^{1},u_{4i+r}^{2})$ and $(u_{4i+r}^{3},u_{4i+r}^{4})$, a single doctor $u_{4i+r}^{5}$, and four new hospitals $h_{4i+r}^{s}$ ($1\leq s\leq 4$), all with capacity 1. The preference lists of the newly added agents corresponding to $y_{4i+r}\in Y$ are shown in Figure 8 and the enforcer gadget is illustrated in Figure 9. In the preference list of $y_{4i+r}$ in $I^{\prime}$, $P(y_{4i+r})$ represents the preference list of $y_{4i+r}$ in $I$.
$\begin{array}[]{rll}(u^{1}_{4i+r},u^{2}_{4i+r}):&(h^{1}_{4i+r},h^{2}_{4i+r}){}{}{}{}&(0\leq i\leq n-1,0\leq r\leq 3)\\ (u^{3}_{4i+r},u^{4}_{4i+r}):&(h^{1}_{4i+r},h^{4}_{4i+r})\succ(h^{3}_{4i+r},h^{2}_{4i+r}){}{}&(0\leq i\leq n-1,0\leq r\leq 3)\\ u^{5}_{4i+r}:&y_{4i+r}\succ h^{1}_{4i+r}{}{}&(0\leq i\leq n-1,0\leq r\leq 3)\\ \\ y_{4i+r}:&P(y_{4i+r})\succ u^{5}_{4i+r}&(0\leq i\leq n-1,0\leq r\leq 3)\vspace{1mm}\\ h^{1}_{4i+r}:&u^{5}_{4i+r}\succ u^{1}_{4i+r}\succ u^{3}_{4i+r}&(0\leq i\leq n-1,0\leq r\leq 3)\vspace{1mm}\\ h^{2}_{4i+r}:&u^{4}_{4i+r}\succ u^{2}_{4i+r}&(0\leq i\leq n-1,0\leq r\leq 3)\vspace{1mm}\\ h^{3}_{4i+r}:&u^{3}_{4i+r}&(0\leq i\leq n-1,0\leq r\leq 3)\vspace{1mm}\\ h^{4}_{4i+r}:&u^{4}_{4i+r}&(0\leq i\leq n-1,0\leq r\leq 3)\vspace{1mm}\\ \end{array}$